Session: May 1997 (session 2)
Frequency: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Onsala (A. Polatidis at Sat May 31 21:11:41 METDST 1997) Success
No known problems

Medicina (M. Bondi at Sat May 31 21:33:13 METDST 1997) Success
No problem

Westerbork (tony foley at Mon Jun 2 13:23:08 METDST 1997) Minor failures
No known problems

Effelsberg (dgraham at Mon Jun 2 16:07:10 METDST 1997) Minor failures
source change to 3c273 took too long

Torun (K.Borkowski at Mon Jun 2 17:05:13 METDST 1997) Minor failures
Part of the first and whole second scan lost due to the telescope control failure.

Jodrell (Alastair Gunn at Thu Jun 5 10:35:02 METDST 1997) Minor failures
Extreme RFI during this experiment. No other known problems.

Cambridge (Alastair Gunn at Thu Jun 5 12:34:24 METDST 1997) Failure
PI provided the schedule for Cambridge late and used incorrect codes line. Snap file and procedure libraries did not match so BBCs all set incorrectly. There was no time to check this before the schedule started and was not noticed until too late.

Urumqi (LI Suqin at Fri Jun 6 04:36:44 METDST 1997) Success
It is cloudy. clock offset =3.0+-0.5us. No problem found.

Noto (C.Trigilio at Mon Jun 9 15:53:51 METDST 1997) Success
No problem reported

Simeiz (A.Volvatch/LIG at Wed Jun 18 08:50:49 METDST 1997) Success
No problem reported Clock offset 0.4 usec. Tapes shipped on 06.06.97.

Calibration (Fredrik T Rantakyro at Mon Jul 21 11:15:45 METDST 1997) Minor failures
The ANTAB file for the experiment can be found on VLBEER in the MAY97 directory with the name C97L2.ANTAB. The file is ready to be used with the AIPS tasks ANTAB/APCAL. Read the ANTAB file for any specific comments regarding this experiment.